If the MPC​ = 0.8, a permanent increase in planned real investment of​ $40 billion will increase real GDP by a total of
Answer #1

Since the given information is;

If the MPC​ = 0.8,

MPS=1-MPC

=1-0.8

=0.2

a permanent increase in planned real investment of​ $40 billion.

Multiplier= 1/ MPS

=1/0.2

=5

Hence a permanent increase in planned real investment of​ $40 billion will increase real GDP by a total of= Multiplier* change in the investment

=5*40 billions

=$200 billions